how do you solve this proportion. x over 3 equals 4 over 6
check for extraneous solutions?

Answer by IWork4Dessert(60)

x/3=4/6
It's not nearly as complicated as that. Since it's basically just division in both parts of the equation(x/3 is just x divided by 3, really, same with 4/6), you can cancel out the bottom parts of your equation. So as the opposite of division is multiplication, multiply both sides by the least common multiple to cancel them all out.
For example:
6 (x/3) = 6(4/6)
Now it's just cross cancellation. Cancel out the three and the six first; as 3 goes into 6 twice, you'll have a 2 left over on the top of the equation. On the other side, 6 goes into 6 once, so nothing's left over and the 4 remains as is.
2x=4
Divide both sides by 2 to get your variable alone(2x is just like 2(4) or whatever else you can plug in there; that's multiplication, so the opposite is division).
x=2
There you go. Hope this helps!